# Variability of X-ray polarization of Cyg X-1

### v1

@@ 3.4 Multiwavelength behavior
Current radio polarization data are not sufficient to trace any signs of orbital or superorbital variability, however, the PAs of all individual datasets are well aligned with the optical and X-ray PAs. The radio PD, on the other hand, seems to be sensitive to the state transitions: nearly 5-fold increase of PD is observed during the transition to the soft state (around Epoch 13). At the same time, the X-ray polarization shows a pronounced drop. Hence, if any connection exists, the radio polarization appears to be anti-correlated with the X-ray polarization.
@@ 4.1 Source of orbital variability
Several processes may induce orbital variations of the observed X-ray polarization in BHXRBs: reflection of X-rays off the companion star (Ahlberg et al., 2024; Rankin et al., 2024) or from the bow shock (V. Ahlberg et al. 2025, in prep.), scattering in the stellar or accretion disk wind (Kallman et al., 2015; Nitindala et al., 2025), or changes in the accretion disk orientation (Bochkarev & Karitskaya, 1983). Below we consider two major classes of models: (i) variations of orientation of the region where the X-ray polarization is produced and (ii) scattering of the incident emission at the intrabinary structure. The models are constrained by the observed criteria imposed by polarization properties: (a) one loop in the -plane per orbital cycle and (b) counter-clockwise rotation (Sect. 3.2 and the increasing angle with the orbital phase in Fig. 7).
@@
The variations of X-ray polarization can be produced by the changing orientation of the site of X-ray production, we can apply the formalism developed for the rotating vector model (Radhakrishnan & Cooke, 1969; Meszaros et al., 1988) to constrain the tilt of the X-ray-emitting inner flow (disk) with respect to the orbital plane. Variations in the PA in this scenario are attributed to changes in the position angle of the projected inner flow axis on the sky. If the disk axis changes its orientation as a function of orbital phase, the PA will vary accordingly. The limited range of PA variations (Fig. 5) corresponds to a small tilt, , which can account for the observed changes. The observed PD variations are caused by the changing inclination. We consider a simplified model in which the PD depends on the cosine of the angle between the disk axis and the observer line of sight, given by . While the actual angular dependence may be more complex, it can be approximated as a linear function of within a narrow angular range. The black solid lines in Fig. 5 (loop in the -plane and sine waves in PD, PA representation) correspond to the best-fit predictions of this model with the tilt , % and fixed orbital inclination .
@@
Changes of the inner disk/flow orientation with respect to the average direction of the orbital axis can be related to tidal forces or precession. In the case of tidal locking, the orientation of the disk is expected to be fixed with respect to the line connecting the compact object and companion star (e.g., always facing the companion star). In this case, the movement of the disk axis on the sky follows the sense of rotation of the compact object around the supergiant (see Fig. 9). This, however, is not observed: the criterion (b) implies that rotation of polarization is counter-clockwise, while the binary system is undergoing clockwise rotation on the sky (; Miller-Jones et al., 2021).
@@
Alternatively, the disk may experience precession in the direction opposite to the direction of orbital motion (retrograde precession). Furthermore, if the precession is synchronous with the orbital period, it should proceed in the binary system with companions of comparable masses, low eccentricity and no known third-body influence. If all these criteria are satisfied, the precession can explain the observed orbital changes of X-ray polarization, however, no known binary system currently fits all of the aforementioned criteria. Hence, we consider this scenario implausible.
@@
In the second class of models (ii), the changing polarization is produced by the scattering (or reflection) on the intrabinary structure. The PD in this case depends on the cosine of the scattering angle, , as (single-scattering case). This relation predicts two PD peaks per orbital period, as is identical for all diametrically opposite points along the orbit. To produce only one peak per orbit and comply with criterion (a), the scattering matter should either be asymmetrically distributed relative to the orbital plane, or the fraction of scattered radiation should vary significantly with the orbital phase. The latter case has been considered in the context of an asymmetric bow shock producing pronounced orbital variability of X-ray polarization in Cyg X-3 (Veledina et al., 2024, and V. Ahlberg et al. in prep.). Polarization in this case can be orthogonal to the orbital axis, hence, it is effectively subtracted from the average polarization of the source, leading to higher net PD at phases when the contribution of the scattering is low. This makes this scenario potentially compliant with criterion (b), however, further quantitative study is needed to draw conclusions on the applicability of this scenario.
@@ 4.2 Long-term trend in polarization
The long-term stability of the X-ray polarization in Cyg X-1 is remarkable and indicates the stability of geometry of the site of polarization production. The statistically significant increase of PD with energy found in both hard and soft spectral states, along with the constant PA across the IXPE energy band (Sect. 3.1) imposes strong constraints on the Faraday rotation effects within the medium responsible for the production of the observed light and any screen between this site and the observer. These constraints translate to upper limits on the large-scale magnetic fields, e.g. a large-scale vertical field G can induce the rotation of PA by across the IXPE band (Barnier & Done, 2024). Our updated estimates on the energy dependence of PA can further refine the constraints on the magnitude of -fields in Cyg X-1.
The statistically significant increase of PD with energy detected in the hard state challenges current models of polarization production. The spectrum is shaped by multiple Compton up-scatterings, with the PD increasing with each successive scattering order (e.g., Poutanen & Svensson, 1996). One way to account for the high observed PD (given a fixed system inclination) in this state is to assume that the IXPE band is dominated by high-order Compton scatterings. This is equivalent to assuming a low energy for the seed photons, such as those originating from synchrotron radiation (e.g., Poutanen & Vurm, 2009; Malzac & Belmont, 2009; Veledina et al., 2013). However, this scenario leads to a suppression of any energy dependence of polarization, since the PD is known to saturate at high (5th) scattering orders (Poutanen & Svensson, 1996; Poutanen et al., 2023). An alternative solution is to retain seed photons from the underlying accretion disc (the slab corona geometry), thus keeping the scattering orders in the IXPE band low to preserve the energy dependence of polarization, and enhance the PD by invoking a bulk matter outflow (Beloborodov & Poutanen, 1999; Poutanen et al., 2023). However, realistic outflow velocities have been found insufficient to reproduce the observed % PD at an inclination for this geometry (Poutanen et al., 2023).
@@
The scenario with the dominant role of returning radiation implies the switch of the dominant polarization production mechanism from Compton scattering in the hard state to reflection of the self-irradiating disk emission in the soft state, while the scenario with hybrid Comptonization implies gradual change of electron temperature, hence, polarization properties, across spectral transition. Our statistical tests (Sect. 3.1) strongly favor a gradual linear change in PD with hardness over the scenario involving two independent polarization mechanisms operating in the soft and hard states.
Interestingly, the increasing PD with energy is expected in both scenarios. The disk emission generally has lower polarization as compared to the reflected (self-irradiated) part that becomes dominant at energies above the disk peak (Schnittman & Krolik, 2009), leading to the natural increase of PD with energy. In the hybrid Comptonization scenario, the increase of PD is attributed to both the low disk intrinsic polarization and the presence of the first Compton scattering order that is polarized in the direction along the disk plane (Poutanen & Svensson, 1996). The dominant role of the disk emission in the IXPE band during the soft state can likewise explain the observed increase of total PD with hardness (Fig. 4).
@@ 5 Summary
We present the results of a comprehensive three-year observational campaign of Cyg X-1 using X-ray, optical, and radio polarimetry. Thirteen IXPE observations conducted between 2022 and 2024 reveal a clear dependence of the X-ray polarization on the spectral state. We find the PD to be approximately twice as high in the hard state () as compared to the soft state (), and stable over time. Furthermore, a model in which PD increases linearly with spectral hardness is favored over one with two independent polarization levels for the different states.
@@ Acknowledgements.
The Imaging X-ray Polarimetry Explorer (IXPE) is a joint US and Italian mission. The US contribution is supported by the National Aeronautics and Space Administration (NASA) and led and managed by its Marshall Space Flight Center (MSFC), with industry partner Ball Aerospace (contract NNM15AA18C). The Italian contribution is supported by the Italian Space Agency (Agenzia Spaziale Italiana, ASI) through contract ASI-OHBI-2022-13-I.0, agreements ASI-INAF-2022-19-HH.0 and ASI-INFN-2017.13-H0, and its Space Science Data Center (SSDC) with agreements ASI-INAF-2022-14-HH.0 and ASI-INFN 2021-43-HH.0, and by the Istituto Nazionale di Astrofisica (INAF) and the Istituto Nazionale di Fisica Nucleare (INFN) in Italy. This research used data products provided by the IXPE Team (MSFC, SSDC, INAF, and INFN) and distributed with additional software tools by the High-Energy Astrophysics Science Archive Research Center (HEASARC), at NASA Goddard Space Flight Center (GSFC). This research has made use of data from the RoboPol programme, a collaboration between Caltech, the University of Crete, IA-FORTH, IUCAA, the MPIfR, and the Nicolaus Copernicus University, which was conducted at Skinakas Observatory in Crete, Greece. We thank the staff of Lord’s Bridge, Cambridge, for the their support in making the AMI observations. This research has been supported by the Finnish Cultural Foundation (VK), the Academy of Finland grant 355672 (AV) and the Magnus Ehrnrooth foundation (APN). MD, JPod, and JS acknowledge the support from the Czech Science Foundation project GACR 21–06825X and the institutional support from the Astronomical Institute RVO:6798581. FM, ADM, and FLM are partially supported by MAECI with grant CN24GR08 “GRBAXP: Guangxi-Rome Bilateral Agreement for X-ray Polarimetry in Astrophysics”. AI acknowledges support from the Royal Society. IL was funded by the European Union ERC-2022-STG-BOOTES-101076343. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. SAT is supported by the Ministry of Science and Higher Education of the Russian Federation grant 075-15-2022-262 (13.MNPMU.21.0003). MB acknowledges the support from GAUK project No. 102323. GM acknowledges financial support from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101107057. POP acknowledges financial support from the Action Thématique PEM from CNRS and from the french spacial agency CNES. AAZ has been supported by the Polish National Science Center grants 2019/35/B/ST9/03944 and 2023/48/Q/ST9/00138.References
@@
- Arnaud (1996) Arnaud, K. A. 1996, in ASP Conf. Ser., Vol. 101, Astronomical Data Analysis Software and Systems V, ed. G. H. Jacoby & J. Barnes (San Francisco: Astron. Soc. Pac.), 17–20
@@
- Ramachandran et al. (2025) Ramachandran, V., Sander, A. A. C., Oskinova, L. M., et al. 2025, A&A, submitted, arXiv:2504.05885

### v2

@@ 3.4 Multiwavelength behavior
Current radio polarization data are not sufficient to trace any signs of orbital or superorbital variability; however, the PAs of all individual datasets are well aligned with the optical and X-ray PAs. The radio PD, on the other hand, seems to be sensitive to the state transitions: nearly five-fold increase in PD is observed during the transition to the soft state (around Epoch 13). At the same time, the X-ray polarization shows a pronounced drop. Hence, if any connection exists, the radio polarization appears to be anti-correlated with the X-ray polarization.
@@ 4.1 Source of orbital variability
Several processes may induce orbital variations of the observed X-ray polarization in BHXRBs: reflection of X-rays off the companion star (Ahlberg et al., 2024; Rankin et al., 2024) or from the bow shock (V. Ahlberg et al., in prep.), scattering in the stellar or accretion disk wind (Kallman et al., 2015; Nitindala et al., 2025), or changes in the accretion disk orientation (Bochkarev & Karitskaya, 1983). Below we consider two major classes of models: (i) variations of orientation of the region where the X-ray polarization is produced and (ii) scattering of the incident emission at the intrabinary structure. The models are constrained by the observed criteria imposed by polarization properties: (a) one loop in the -plane per orbital cycle and (b) counter-clockwise rotation (Sect. 3.2 and the increasing angle with the orbital phase in Fig. 7).
@@
The variations of X-ray polarization can be produced by the changing orientation of the site of X-ray production, we can apply the formalism developed for the rotating vector model (Radhakrishnan & Cooke, 1969; Meszaros et al., 1988) to constrain the tilt of the X-ray-emitting inner flow (disk) with respect to the orbital plane. Variations in the PA in this scenario are attributed to changes in the position angle of the projected inner flow axis on the sky. If the disk axis changes its orientation as a function of orbital phase, the PA will vary accordingly. The limited range of PA variations (Fig. 5) corresponds to a small tilt, , which can account for the observed changes. The observed PD variations are caused by the changing inclination. We consider a simplified model in which the PD depends on the cosine of the angle between the disk axis and the observer line of sight, given by . While the actual angular dependence may be more complex, it can be approximated as a linear function of within a narrow angular range. The solid black lines in Fig. 5 (loop in the -plane and sine waves in PD, PA representation) correspond to the best-fit predictions of this model with the tilt , % and fixed orbital inclination .
@@
Changes of the inner disk or flow orientation with respect to the average direction of the orbital axis can be related to tidal forces or precession. In the case of tidal locking, the orientation of the disk is expected to be fixed with respect to the line connecting the compact object and companion star (e.g., always facing the companion star). In this case, the movement of the disk axis on the sky follows the sense of rotation of the compact object around the supergiant (see Fig. 9). This, however, is not observed: the criterion (b) implies that rotation of polarization is counter-clockwise, while the binary system is undergoing clockwise rotation on the sky (; Miller-Jones et al., 2021).
@@
Alternatively, the disk may experience precession in the direction opposite to the direction of orbital motion (retrograde precession). Furthermore, if the precession is synchronous with the orbital period, it should proceed in the binary system with companions of comparable masses, low eccentricity and no known third-body influence. If all these criteria are satisfied, the precession can explain the observed orbital changes of X-ray polarization; however, no known binary system currently fits all of the aforementioned criteria. Hence, we consider this scenario implausible.
@@
In the second class of models (ii), the changing polarization is produced by the scattering (or reflection) on the intrabinary structure. The PD in this case depends on the cosine of the scattering angle, , as (single-scattering case, Chandrasekhar 1960). This relation predicts two PD peaks per orbital period, as is identical for all diametrically opposite points along the orbit. To produce only one peak per orbit and comply with criterion (a), the scattering matter should either be asymmetrically distributed relative to the orbital plane, or the fraction of scattered radiation should vary significantly with the orbital phase. The latter case has been considered in the context of an asymmetric bow shock producing pronounced orbital variability of X-ray polarization in Cyg X-3 (Veledina et al., 2024, and V. Ahlberg et al. in prep.). Polarization in this case can be orthogonal to the orbital axis, hence, it is effectively subtracted from the average polarization of the source, leading to higher net PD at phases when the contribution of the scattering is low. This makes this scenario potentially compliant with criterion (b); however, further quantitative study is needed to draw conclusions on the applicability of this scenario.
@@ 4.2 Long-term trend in polarization
The long-term stability of the X-ray polarization in Cyg X-1 is remarkable and indicates the stability of geometry of the site of polarization production. The statistically significant increase in PD with energy found in both hard and soft spectral states, along with the constant PA across the IXPE energy band (Sect. 3.1) imposes strong constraints on the Faraday rotation effects within the medium responsible for the production of the observed light and any screen between this site and the observer. These constraints translate to upper limits on the large-scale magnetic fields, for instance, a large-scale vertical field G can induce the rotation of PA by across the IXPE band (Barnier & Done, 2024). Our updated estimates on the energy dependence of PA can further refine the constraints on the magnitude of -fields in Cyg X-1.
The statistically significant increase in PD with energy detected in the hard state challenges current models of polarization production. The spectrum is shaped by multiple Compton up-scatterings, with the PD increasing with each successive scattering order (e.g., Poutanen & Svensson, 1996). One way to account for the high observed PD (given a fixed system inclination) in this state is to assume that the IXPE band is dominated by high-order Compton scatterings. This is equivalent to assuming a low energy for the seed photons, such as those originating from synchrotron radiation (e.g., Poutanen & Vurm, 2009; Malzac & Belmont, 2009; Veledina et al., 2013). However, this scenario leads to a suppression of any energy dependence of polarization, since the PD is known to saturate at high (5th) scattering orders (Poutanen & Svensson, 1996; Poutanen et al., 2023). An alternative solution is to retain seed photons from the underlying accretion disc (the slab corona geometry), thus keeping the scattering orders in the IXPE band low to preserve the energy dependence of polarization, and enhance the PD by invoking a bulk matter outflow (Beloborodov & Poutanen, 1999; Poutanen et al., 2023). However, realistic outflow velocities have been found insufficient to reproduce the observed % PD at an inclination for this geometry (Poutanen et al., 2023).
@@
Interestingly, the increasing PD with energy is expected in both scenarios. The disk emission generally has lower polarization as compared to the reflected (self-irradiated) part that becomes dominant at energies above the disk peak (Schnittman & Krolik, 2009), leading to the natural increase in PD with energy. In the hybrid Comptonization scenario, the increase in PD is attributed to both the low disk intrinsic polarization and the presence of the first Compton scattering order that is polarized in the direction along the disk plane (Poutanen & Svensson, 1996). The dominant role of the disk emission in the IXPE band during the soft state can likewise explain the observed increase of total PD with hardness (Fig. 4).
@@ 5 Summary
We present the results of a comprehensive three-year observational campaign of Cyg X-1 using X-ray, optical, and radio polarimetry. Thirteen IXPE observations conducted between 2022 and 2024 reveal a clear dependence of the X-ray polarization on the spectral state. We find the PD to be approximately twice as high in the hard state () as compared to the soft state (), and stable over time.
@@ Acknowledgements.
The Imaging X-ray Polarimetry Explorer (IXPE) is a joint US and Italian mission. The US contribution is supported by the National Aeronautics and Space Administration (NASA) and led and managed by its Marshall Space Flight Center (MSFC), with industry partner Ball Aerospace (contract NNM15AA18C). The Italian contribution is supported by the Italian Space Agency (Agenzia Spaziale Italiana, ASI) through contract ASI-OHBI-2022-13-I.0, agreements ASI-INAF-2022-19-HH.0 and ASI-INFN-2017.13-H0, and its Space Science Data Center (SSDC) with agreements ASI-INAF-2022-14-HH.0 and ASI-INFN 2021-43-HH.0, and by the Istituto Nazionale di Astrofisica (INAF) and the Istituto Nazionale di Fisica Nucleare (INFN) in Italy. This research used data products provided by the IXPE Team (MSFC, SSDC, INAF, and INFN) and distributed with additional software tools by the High-Energy Astrophysics Science Archive Research Center (HEASARC), at NASA Goddard Space Flight Center (GSFC). This research has made use of data from the RoboPol programme, a collaboration between Caltech, the University of Crete, IA-FORTH, IUCAA, the MPIfR, and the Nicolaus Copernicus University, which was conducted at Skinakas Observatory in Crete, Greece. We thank the staff of Lord’s Bridge, Cambridge, for their support in making the AMI observations. This research has been supported by the Finnish Cultural Foundation (VK), the Academy of Finland grant 355672 (AV) and the Magnus Ehrnrooth foundation (APN). MD, JPod, and JS acknowledge the support from the Czech Science Foundation project GACR 21–06825X and the institutional support from the Astronomical Institute RVO:6798581. FM, ADM, and FLM are partially supported by MAECI with grant CN24GR08 “GRBAXP: Guangxi-Rome Bilateral Agreement for X-ray Polarimetry in Astrophysics”. AI acknowledges support from the Royal Society. IL was funded by the European Union ERC-2022-STG-BOOTES-101076343. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. SAT is supported by the Ministry of Science and Higher Education of the Russian Federation grant 075-15-2022-262 (13.MNPMU.21.0003). MB acknowledges the support from GAUK project No. 102323. GM acknowledges financial support from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101107057. POP acknowledges financial support from the Action Thématique PEM from CNRS and from the french spacial agency CNES. AAZ has been supported by the Polish National Science Center grants 2019/35/B/ST9/03944 and 2023/48/Q/ST9/00138.References
@@
- Ramachandran et al. (2025) Ramachandran, V., Sander, A. A. C., Oskinova, L. M., et al. 2025, A&A, 698, A37